Two numbers when divided by 17, leave remainders 13 and 11 respectively.If the sum of those numbers is divided by 17, then what will be the remainder?

Grade:9

1 Answers

Ravi
9 years ago
The 2 no.s can be written in the form of 17n+13 and 17m+11.n.m belong to integers.
Sum =17(m+n)+24
This number when divided by 17 would leave a remainder 7 as

Sum= 17(m+n)+17+7
= 17(m+n+1)+7
.........................................................................................................................................